9. Problems and Applications Q3 Pat and Kris are roommates. They spend most of their time studying (of course), but they leave some time for their favorite activities: making pizza and brewing root beer. Pat takes 4 hours to brew a gallon of root beer and 2 hours to make a pizza. Kris takes 6 hours to brew a gallon of root beer and 4 hours to make a pizza. Pats opportunity cost of making a pizza is 1/2 gallon of root beer, and Kriss opportunity cost of making a pizza is 2/3 gallon of root beer Pat has an absolute advantage in making pizza, and Pat has a comparative advantage in making pizza. If Pat and Kris trade foods with each other,will trade away pizza in exchange for root beer. The price of pizza can be expressed in terms of gallons of root beer. The highest price at which pizza can be traded that would make both roommates better off is of root beer, and the lowest price that makes both roommates better off is of root beer per pizza.
tand Kris are roommates. They spend most of their time studying (of course), but they leave some time for d brewing root beer. Pat takes 4 hours to brew a gallon of root beer and 2 hours to make a pizza. Kris takes d 4 hours to make a pizza. ts opportunity cost of making a pizza is ot beer. 1/2 gallon of root beer, and Kriss opportunity cost of making 1/2 gallon Pat advantage in making pizza, and Pathas a comparative advantage in making pi 2 gallons Pat and K2/3 gallon ith each other,will trade away pizza in exchange for root beer 1 1/2 gallons The highest price at which pizza can be traded th pressed in terms of gallons of root beer. e price of etter off is of root beer, and the lowest price that makes both roommates better off is Grade It Now Capyright NoticesTerms of Use
ns Q3 They spend most of their time studying (of course), but they leave some time for their favorite activities: making kes 4 hours to brew a gallon of root beer and 2 hours to make a pizza. Kris takes 6 hours to brew a gallon of root ing a pizza is _1/2 gallon of root beer, and Kriss opportunity cost of making a pizza is 2/3 gallona 1/2 gallon 2 gallons 2/3 gallon 1 1/2 gallons advantage in making pizza, and_Pat has a comparative advantage in m ith each other, will trade away pizza in exchange for root beer. ressed in terms of gallons of root beer. The highest price at which pizza can be d make both roommates of root beer per pizza. of root beer, and the lowest price that makes both roommates better off is Grade It Now Save & Continue Continue without saving
Answer #1

Pat's opportunity cost of making a pizza is = 2/4 = 1/2 gallon of root beer

Kris's opportunity cost of making a pizza = 4/6 = 2/3 gallon of root beer

Pat has an absolute advantage in making pizza (lesser time taken) and Pat has a comparative advantage in making pizza (lower opportunity cost)

Pat will trade away pizza in exchange for root beer. (Pat has a comparative advantage in making Pizza)

Highest price of pizza = 2/3 gallon of root beer and lowest price of pizza = 1/2 gallon of root beer.